Water at 340 K and a flow rate of 5 kg/s enters a black, thin-walled tube, which passes through a large furnace whose walls aWrite an expression for the linearized radiation coefficient corresponding to radiation exchange between the outer surface of the pipe and the furnace walls. Determine how to calculate this coefficient if the surface temperature of the tube is represented by the arithmetic mean of its inlet and outlet values. Use these expressions to determine the outlet temperature of the water, T m , o , in K.
Answer #1

Tube, D = 0.25 m L = 8 m, E = 1 Air T = 700 K Water m = 5 kgis 1 T. = 300 K Furnace. Tu = 700 KKNOWN: Water at prescribed temperature and flow rate enters a 0.25 m diameter, black thin-walled tube of 8- m length, which pIm =(Tm,i +Imo)/2 1 Rtot = Revi+ 1/Revo +1/R rad Im - 7 It - Tfur Revi 1/Rev.o +1/Rrad The thermal resistances, with As =PL =Note:plzzz don't give dislike...plzzz comment if you have any problem i will try to solve your problem...plzzz give thumbs up i am in need.....